OpenStudy (steve816):

Quick really easy chemistry question! The element barium forms ions with a +2 charge. How many electrons do these ions contain?

OpenStudy (jadedry):

I know it's confusing (to me at least haha) but positive charges are created when an atom /loses/ electrons. So, the normal number of electrons in barium - (+2) = 56 - 2 = 54.

OpenStudy (steve816):

Thank you!

OpenStudy (michele_laino):

In order to form a +2e ion, barium has to lose 2 electrons, so it will remain with 56-2=54 electrons